OpenStudy (anonymous):

Given f(x) = sqrt(1+x) . Use the definition of the derivative to find f’(x). No points will be given for other method of solutions. Find the equation of the tangent line at the point (3, 2).

OpenStudy (danjs):

limit definition?

OpenStudy (danjs):

as h goes to zero

OpenStudy (danjs):

\[\frac{ f(a+h)-f(a) }{ h }\]

OpenStudy (anonymous):

I got 1/(sqrt1+x)(sqrt1+x)

OpenStudy (anonymous):

\[\frac{ 1 }({\sqrt{1+x})(\sqrt{1+x}) }\]

OpenStudy (anonymous):

that didn't come out as I wanted...

OpenStudy (danjs):

\[\lim_{h \rightarrow 0}\frac{ f(x+h) - f(x) }{ h }\] put in f(x+h) and f(x) into it

OpenStudy (anonymous):